Project: Pueblo Viejo
Report: BGC Weekly Geotechnical Summary
Date: May 25, 2012 to May 31, 2012
The following is a brief summary of the earthwork activities noted by BGC Engineering Inc. (BGC) for the
Pueblo Viejo Dominican Corporation (PVDC) Pueblo Viejo project during the period May 25 to May 31, 2012.
Weather predominately hot and dry with no rain; no red alerts.
1.0
GENERAL MANAGEMENT AND ADMINISTRATIVE ITEMS:

K-122 LL Dam Daily (4 PM) construction co-ordination meetings continued through this period (including HYT,
PVDC Construction, Concremat, Epsa Labco, and BGC).

Concremat chair the weekly K-122 (LLSD) co-ordination meeting at 5 PM on Thursday and the K122 contract
meeting on Wednesdays at 2:30 PM. PVDC-CM chair meeting for ARD Dams 1 and 3 10:30 AM on Wed.
THIS WEEK’S ACTIVITIES
GENERAL PROGRESS AND NOTES:
2.0 EL LLAGAL TAILINGS STORAGE FACILITY:
LL Coffer Dam, Starter Dam and Seepage Recovery Dam Construction






Trimming and foundation preparation (LPF zone)
o Right Abutment to el. 175
o Left Abutment to 171 m.
o Trimming RA DS continued.
Foundation approvals:
o Right Abut to 173
o Left Abut to 171
Fill placement (LPF) this week was
approx.55,300m3 (T&S Data).
Fill Type
Average Elevation
LPF
170
Fine Filter
171
Coarse filter
171
Transition Rock Fill
171
D/S Rock fill
171
U/S Rock Fill
171
U/S Rock Fill shell
171
QA QC summary of activities at LL Starter Dam
for the week are included at the end of this report.
NCRs.
o New - no new NCR’s this week
o Remains open
 NCR 12 -12 – Delay of Rock Fill
density test
BGC working with
EPSA – test started this week
 NCR 12 – 15 – Timely Reporting
of QC Test Results. BGC request
update from Epsa Labco and
Concremat on the request for
additional staff and facilities made
by Epsa Labco.
o Closed – 0






Water level in the Cofferdam Reservoir was at elev.
139.9 m on May 31. This is a decrease of 2.3 m from
last week. The chart attached to this report shows
that approx.1.5 million cubic meters of runoff water is
currently stored in the reservoir. Total reservoir
capacity is approx. 17 million cubic meters at the
current dam crest of approx.170 m.
Six @ 100 to 125 HP submersible pumps were
operated for 6 days this week
Permanent water reclaim system ran for 7 days this
week discharging with 2 of 3 pumps at a rate of
1150 m3/hr each.
Two siphons are available for emergency use.
Summary of Pumping Activity
PUMP TYPE
M3/hr./ea
Days/week
PVDC Barge (1)
145
6
PVDC Barge (1)
145
6
HYT Sub.
255
6
PVDC Reclaim (2)
1150
7
PVDC Diesel (1)
N/A
(4)
Regular monitoring of seepage at the D/S toe
continued. Flow measurements are being tabulated
for comparison and trend. Rainfall runoff is adding to
discharge. Observations will be maintained.
INDRHI inspection on May 25
Rock fill density test was started at the end of this
report period
LL Seepage Recovery Dam






The grouting program was completed on
May 19th
Fill placement started this week.
2.0 m. of LPF have been placed in the U/S
core to aprox... elev. 208m.
1.5 m the D/S blanket was also placed
Foundation prep and water control were
ongoing.

The INDRHI permit to place fill was given this
week
Field Memo #20 was issued which details the
excavation of the LA alluvial soils after the waste
dump above is removed.
Low Permeability Fill Borrow





LPF material was extracted from borrow areas “H,
and L” this week.
BGC-QA staff visited all borrows currently being
utilized or containing stockpiled material (H, H1, I,
and L). Epsa Labco QC staff also inspecting
borrow.
Borrow “I” continued to produce low plastic LPF
along with poor mixing Borrow “I” was shut
down on May 28
Isolated cases of high moisture content in the LPF
borrow are still being detected but are not a major
problem to production or quality control at this
time.
With the hot dry windy weather drying of wet
LPF is done quickly

PVDC-CM stockpile volume of LPF in Borrow Area
“H” and “L” is approx. 35,100 m3; as reported by
PVDC this week.
Filters and Rockfill






Delivery of sand and gravel from Rio Yuna for the
production of fine and coarse filter continued this
week.
BGC continue weekly inspection of Hermanos
Yarull aggregate extraction source at the Rio
Yuna. No significant changes in the “raw granular
product” delivered to El Llagal stockpiles.
HYT continue production of fine and coarse filter
at El Llagal #2 and Los Bananos.
BGC visit the stockpiles on several occasions
during the reporting period with Epsa-Labco QC.
The “off-site” sources of Cabrera, and Diaz in
Cotui continue to deliver fine and coarse filter.
Pressa Issa with help from EPSA Labco have
been working at producing an in spec product.
Filter Source
Llagal, Los
Bananos
Cabrera
Diaz
Fine (m3)
Coarse (m3)
6300*
10100*
5400**
748**
8800**
* As reported by HYT.
** As reported by PVDC
*** Estimate


PVDC’s stockpile report of May 31th, 2012 indicated
the following:
o Rock fill
52,300m3
o Fine filter
4,280m3
o Coarse filter
57,500m3
o Raw river rock
35,300m3
o LPF
35,100m3
T&S/ Concremat report of June 3rd indicates that:
115,300 m3 of fine filter and 104,800 of coarse filter
have been placed in the LL Starter Dam.
3.0 ARD DAM 1 AND BASIN AND SEEPAGE RECOVERY DAM (SRD)








ARD Dam 1 - No work related activities.
Dam 1 Basin – Two new diesel pumps (2300
gal/min) operating. The water level has dropped
by approximately another 2 meters.
It is planned to draw the pond down completely
and make the final repairs to the liner.
PVDC completed the construction of the French
drain above the inlet channel at the North side of
the basin. Some water is flowing from the drain
toward the SRD reservoir.
Dam 1 Seepage Recovery Dam - HYT placed
material in all zones of the Dam. BGC worked
with PVDC-CM and HYT QC to achieve
compliance with the drawings and technical
specifications.
The elevation of the fill placement is as follows:
Fill Type
Average Elevation
masl
LPF
143
Fine Filter
143
Coarse Filter
143
Rock Fill
144
Fluor continues working on the installation of the
permanent SRD pumping system.
Dam 1 Spillway - No work related activities

HYT submitted an update on the Deficiency
List, including: ( no major change this
week)




Geomembrane deficiency repairs continue
This list is under review by BGC. (approx.
70% of the original items have been
addressed by the contractor )
Fluor have not started on the work for the
energy dissipation structures at the 3
culvert inlets
No work has been started to repair the HR2 ditch.
4.0 ARD DAM 3 AND SEEPAGE RECOVERY DAM (SRD)
ARD Dam 3:





The work has been going well with good
progress; the hot dry weather has helped.
The left abutment is still lagging and has needed
more resources. PVDC imported 20 more
labourers to work this area where a depression
has formed.
HYT continued placing rock fill material on the
upstream and downstream zones of the dam.
Summary of fill elevations follows:
Fill Type
Average
Elevation masl
LPF
295.5
Fine and Coarse filters
295.6
upstream
Rock fill upstream
296
Fine and Coarse filters
295.5
downstream
Rock fill downstream
296.1
Dam 3 Cofferdam:
 One pump (2300 gal/min) required minimally during
the day. The water level remains at a low level in the
reservoir.
Dam 3 SRD:
 No work related activities.
Dam 3 Spillway:
HYT continued working on the concrete road crossing /
control structure of the emergency spillway. The center
portion was poured this week. One lane is complete.
NCRs.
o Closed – 3 (Dam 3)
 NCR 12-16 – LPF Fill to FF fill
Sequence.
 NCR 12-17 – Foundation
Preparation and contact LPF.
 NCR 12-18 – Poor surface water
Management.
5.0 UM-1 & UM-2; HONDO DAM

BGC continue to monitor the existing slope indicators at the Hondo Dam.

The trees and scrub vegetation were removed from the dam this week.
Additional instrumentation will be installed at the Hondo Dam

6.0 FLUOR PLANT SITE EARTH WORKS



BGC assessing the landslide below Lime Kilns. Seepage path into slide area was observed.
BGC were requested to assess the slope near the Mill Office on Middle bench.
BGC were requested to assess an alternate location for the Bio-Mass Pond. The proposed location is
north of the current site and is in a localized area of steeper topography. The BGC assessment is near
completion.
7.0 BGC SITE MANPOWER (WEEK ENDING 31, MAY, 2012)
Capacity
Construction
Engineering
Contractor
Local
support
Review
Engineers
Total
Personnel
Support BGC
Equipment Operators Forman Staff
Engineers
7
Man Man
Days Hours Notes
49
588
2*
11
65
520
2
18
114
1108
31 MAY 2012 VIEW OF LL STARTER DAM LOOKING TO THE EAST.
*
Secretary
SEE NEXT PAGES FOR QA /QC REPORTS
AND STORAGE CAPACITY OF LLSD
QA /QC REPORT FOR LLSD ( QC BY EPSA LABCO)
QA /QC REPORT FOR ARD DAM 1 AND DAM 3 (QC BY HYT)
RESERVOIR STORAGE CAPACITY/ STORED WATER GRAPH -- MAY 31,2012